Please show work.
Consider the quadratic equations:
x^2-2=0
Solve quadratic equation above by
extracting square roots
completing the square
by applying the formula

Answer by Boreal(15235)

x^2=2
x= +/- sqrt (2)
-----------------
x^2-2=0
(x-sqrt(2))^2=0
x=+/- sqrt (2)
--------------
x=(1/2)(0+/- sqrt (0+8)
This is (1/2)*+/- 2 sqrt(2)
That is +/-sqrt (2)